Property sold for $330,000 with 10% down, financed at 7.5% interest for 20 years (monthly payments). The property has a gross annual income of $42,000. Vacancy collection loss is 8% and expenses are $8,000 per year. What is the equity dividend rate indicated by this sale?
Answer #1

Property Price = $ 330000, Interest Rate = 7.5 % and Loan Tenure = 20 years pr (20 x 12) = 240 months with monthly payments, Downpayment : 10 %

Down Payment = Cash Investment = 0.1 x 330000 = $ 33000 and Borrowing = 330000 - 33000 = $ 297000

Monthly Interest Rate = 7.5/12 = 0.625 %

Let the monthly repayments (monthly debt servicing) be $ K

Therefore, 297000 = K x (1/0.00625) x [1-{1/(1.00625)^(240)}]

297000 = K x 124.132

K = 297000 / 124.132 = $ 2392.61

Annual Debt Servicing = 12 x K = 12 x 2392.61 = $ 28711.3

Gross Annual Income = $ 42000

Less: Vacany and Collection Loss = 8 % of Gross Income = 0.08 x 42000 = $ 3360

Less: Operating Expenses = $ 8000

Net Operating Income (NOI) = 30640

Less: Annual Debt Servicing = $ 28711.3

Pre-Tax Cash Flow (PTCF) = $ 1928.7

Equity Dividend Rate = PTCF / Down Payment = 1928.7 / 33000 = 0.05845 or 5.845 % ~ 5.84 %